OpenStudy (anonymous):

Find all polar coordinates of point P = (2, 14°).

OpenStudy (amorfide):

is that not already polar coordinate form

OpenStudy (welshfella):

do you know what polar co-ordinates look like?

OpenStudy (anonymous):

No and no.

OpenStudy (anonymous):

Wait..yes and no.

OpenStudy (welshfella):

no (2,14) are Cartesian coordinates

OpenStudy (anonymous):

It is but you have to find all ones equivalent? I think?

OpenStudy (amorfide):

thats a 14 degrees though

OpenStudy (welshfella):

oh - sorry - you are right

OpenStudy (amorfide):

well for polar coordinates, to find them all you usually write it as \[(r,\theta + 2\pi(n))\]

OpenStudy (amorfide):

r=2 theta=14

OpenStudy (amorfide):

you already have polar coordinates so the only thing left would be to write it as I shown above using (r, theta + 2pin) where theta is in radians and n is any integer or (-r, theta) is the same as the plot of \[(r, \theta \pm \pi)\]

OpenStudy (amorfide):

not sure what else to answer to your question sorry

OpenStudy (ybarrap):

We need to covert Cartesian to polar:  Using Euler's Identity: $$ e^{i\theta}=\cos\theta+i\sin \theta $$ In polar coordinates: $$ \large P=re^{i\theta}\\ \theta=\tan^{-1}{14\over2}=\tan^{-1}7=81.9~\text{degrees}\\ r=\sqrt{2^2+14^2}=14.14\\ \implies P=14.14e^{i81.9}\\ $$ Where \(r\) is the distance from the origin to P and \(\theta\) is the angle between the x-axis and segment \(r\) .

OpenStudy (ybarrap):

 see- http://en.wikipedia.org/wiki/Polar_coordinate_system#Complex_numbers